Given the demand function Qd = 40-2p and the supply function 2P-Q = 20. Assume the government impose at a tax of t per unit on quantity supplied and the producer adjust the supply function to include the tax.

a.      Find the tax rate which will maximize the tax revenue?

b.     Find the maximum revenue which can be obtained from taxation?


1
Expert's answer
2022-01-29T11:06:01-0500

A) The Tax Rate At Which Tax Revenu Is Maximum is when t = (Pd - Ps) = Q.

(20 - 0.5Q - 10 - 0.5Q) = Q,

Q = t = 5.


B) The Maximum Revenue Which Can Be Obtained From A Function is:

TR = 5×5 = 25.